Question 3 Magic Corporation (MC) produces special carbon fibre basketball hoops. These hoops are produced in batches, where machines and moulds have to be set up for every batch Setup costs are batch-level costs because they are associated with batches rather than individual units of products. A separate Setup Department is responsible for setting up machines and moulds Setup overhead costs consist of some costs that are variable and some costs that are fixed with respect to the number of setup-hours. The following information pertains to March 2016 Static-budget Amounts Actual Amounts Basketball hoops produced and sold Batch size (number of units per batch) Setup-hours per batch Variable overhead cost per setup hour Total fixed setup overhead costs 60,000 400 10 $20 $45,000 56,000 500 $18 $42,000 Required: (a)Calculate the efficiency variance for variable setup overhead costs (4 marks) (b)Calculate the spending variance for variable setup overhead costs (4 marks) (c) Explain why the spending variance for variable setup overhead costs is the appropriate variance for evaluating the manager in charge of the Setup (2 marks) (d) Calculate the flexible-budget variance for variable setup overhead costs. Write a 6 marks (e)Calculate the spending variance for fixed setup overhead costs. Explain how this (3 marks) Department. journal entry to record the variable setup overhead cost for March 2016 variance can be useful to MC for controlling production costs (f)Calculate the production-volume variance for fixed setup overhead costs (5 marks)
Answer #1

a) Variable efficiency overhead variance: (Standard hours- Actual hours)* Standard Rate

=(500*10-500*8)*20=(5000-4000)*20=1000*20=20000F

b) Variable spending overhead variance: (Standard Rate- Actual Rate)* Actual hours

=(20-18)*4000 = 2*4000 = 8000F

c) Spending variance is appropriate variance for evaluating the manager in charge of set up department because paying a higher / lower average actual overhead price per unit of the activity base than the standard price allowed per unit of the activity base. and. larger / smaller waste and shrinkage associated with the resources involved than expected.

d) Journal entry to record variable overhead set up costs is

Finished goods inventory A/c Dr 28000

To variable overhead eficiency variance 20000

To Variable overhead rate variance 8000

e) Calculation of Expenditure variance for fixed set up overhead costs

Fixed Overhead Expenditure variance= Budgeted hours * Standard rate - Actual hours * Actual rate

=(400*112.5 - 500* 84)= 3000F

This helps the MC for controlling costs

  • Planned business expansion, which was anticipated to cause a stepped increase in fixed overheads, not being undertaken during the period.
  • Cost rationalization measures carried out during the period aimed at reducing fixed overheads by elimination of inefficiencies

f) Fixed overhead volume variance is Standard hours* Standard rate - budgeted hours* Standard rate

=48000-48000 = 0